UT 20231005: ECD - Chiller Maintenance



Night Info

For the night of: Wed, 2023 Oct 4 ---> Thu, 2023 Oct 5
Local midnight = 2023 Oct 5,  7 hr UT, or JD 2460222.792
Local Mean Sidereal Time at midnight =   0 35 15.4

Sunset ( 1000 m horizon):   18 05 MST; Sunrise:   6 11 MST
Evening twilight:  19 22 MST;  LMST at evening twilight:  19 57
Morning twilight:   4 55 MST;  LMST at morning twilight:   5 31
12-degr twilight: 18 53 MST -->  5 23 MST; night center:   0 08 MST

Moonrise:  21 42 MST   
Moon at civil midnight: illuminated fraction 0.620
1.3 days before last quarter, RA and dec:   5 50 24, +27 42.0

The sun is down for 12.1 hr;  9.5 hr from eve->morn 18 deg twilight.
 2.3 dark hours after end of twilight and before moonrise.

Weather:


Staff

  • OSA: David Gonzalez Huerta
  • ISA/SO: Jenny Power
  • ECD Coordinator: Jenny Power


Summary

20231005 Night Log

Critical Note

Chiller work will leave us operating at 50% capacity. Based on heatload assessments done in advance, we will operate only if the 47F is controlled below alarm setpoints. To help reduce load, mirror ventilation will not be run. HBS will be stopped if temperatures get above threshold to ensure the LT chiller (CH0402) that cools ICS on high discharge pressure from 47F getting to warm.

Night Plan

TCS 2023B patch

TCS 2023B will be updated in advance of ECD night with a patch. Details pending.

AZCAM Server update

Testing new AZCAM server update - hotspot and shack-hartmann sensor imaging on new software of azcam. Details pending

FBG focal station checkout
  1. AGW1/2 Alignment
  2. M3 Alignment LFBG,RFBG
  3. Field Aberration

Pointing Model Implementation/Verification

Pending analysis of data from UT20230923

MODS Alignment Verification
  1. Blind Offset test*
    • StoneA or G20p25, execute longslit blind offset acquisition.
  2. Rotator Zero Points
    • Scripts and MMS files are located: /home/lbto/Service_Observing/Restart_2023/MODS/RZP
    • These 3 masks cover the night. Try to align 1 more masks. ~15 min each.
      • One mask was not perfectly aligned. retry:
      • 00p25
      • 04p25
  3. Photometric Standards
    • Available first half: Need to check nimgs in scripts (times based on scripts in Restart_2022).
      • M71 : twilight - 8UT: ~30 min (21 for the script and the extra for slew+collim)
        • dual (8 min)
        • red (7 min)
        • blue (6 min)
      • M92: twilight - 6 UT ~40 min
        • dual (12 min)
        • red (12 min)
        • blue (10 min)
      • M13: twilight - 5 UT
  4. Spectrophotometric Standards
    • BD+33 until 04:30UT, BD+28 until 10UT, Feige110 6-10UT, G191-B2B from 9:30UT to twilight. Try to get Feige110 and/or G191-B2B which means standards are 2nd half of night.
    • If Clear, 2-3 spectra/mode
      For 3 spectra/mode, this will take 1hr/star
    • If not clear, 1 spectrum/mode
      One spectrum/mode will take 36 min
Repoint Checks and Tests

Doug wants to test some flags and methods for the repoint task with AO.
Slope Zero Software Testing
  • Use the LBTI WFS to test the Slope zero software patch
  • To be run in parallel with repoint task.

Closed Dome Plan

  • Generate and test low mode LBTI reconstructor for SHARK-VIS

Night log

-- JenniferPower - 04 Oct 2023
Topic revision: r5 - 16 Oct 2023, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback